OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = -7 3x - y = -5 3x - y = 5 x + 3y = -5

OpenStudy (anonymous):

Gradient = (10 - 1) / (1 - (-2)) = 9 / 3 = 3 Using (-2, 1): y = 1 + 3(x - (-2)) = 1 + 3x + 6 = 3x + 7 3x - y = -7 (answer)

OpenStudy (anonymous):

I don't understand how to do the equation, but thank you very much fro helping me and explaining how you got to the answer!:D